# Prior-based Noisy Text Data Filtering: Fast and Strong Alternative For Perplexity

### v1

@@ Abstract
As large language models (LLMs) are pretrained on massive web corpora, careful selection of data becomes essential to ensure effective and efficient learning. While perplexity (PPL)-based filtering has shown strong performance, it suffers from drawbacks: substantial time costs and inherent unreliability of the model when handling noisy or out-of-distribution samples. In this work, we propose a simple yet powerful alternative: a prior-based data filtering method that estimates token priors using corpus-level term frequency statistics, inspired by linguistic insights on word roles and lexical density. Our approach filters documents based on the mean and standard deviation of token priors, serving as a fast proxy to PPL while requiring no model inference. Despite its simplicity, the prior-based filter achieves the highest average performance across 20 downstream benchmarks, while reducing time cost by over 1000× compared to PPL-based filtering. We further demonstrate its applicability to symbolic languages such as code and math, and its dynamic adaptability to multilingual corpora without supervision. The code is available online (https://github.com/ybseo-ac/prior_filter).
@@ 1 Introduction
The prior-based filter exhibits intriguing and practical properties, which we demonstrate empirically. (1) The linguistic principles underlying the term-frequency hold not only for English but also for other natural languages (e.g., Chinese and French), even for symbolic languages (e.g., code, mathematics). (2) Only a small amount of Chinese data mixed into an English corpus may be noise and models can not learn patterns from it; however, as its amount increases, it becomes learnable by models. The prior-based filter is capable of automatically capturing this transition of learnability.
@@ 3.4.1 Prior-based filter approximates PPL-based filter
Figure 3: Extreme outlier samples selected based on three criteria, ensuring that each sample comes from a distinct criterion: PPL, $\mu_{d}$, and $\sigma_{d}$.
@@
3.4.2 $\mu_{\mathtt{d}}$ reflects language learnability in multi-lingual corpora
The prior mean value has the property of dynamically reflecting the learnability of a data cluster (i.e., language type), especially when multiple clusters with distinct characteristics are mixed. For example, consider a corpus primarily composed of English data with a small portion of Chinese data included. While the Chinese samples may contain meaningful content, if their quantity is too small, the model will fail to learn the language. In this case, Chinese data is no more than noise. However, once the volume of Chinese data increases sufficiently, the model becomes able to interpret the language, making it learnable and meaningful data.
@@ 3.4.3 Fast, scalable filtering using subsampled priors
Remarkably, the already minimal computation time of the prior-based filter can be further reduced. For a 6B-token corpus, the entire process takes about 35 minutes on 40 CPUs (Intel Xeon Silver 4210R @ 2.40GHz), which consists of two stages: assessing the token prior, and computing $\mu_{\mathtt{d}}$ and $\sigma_{d}$. Among these, the most time-consuming step is the token prior assessing phase, which alone takes around 30 minutes.
@@ 4.1 Experiment on natural language corpus and general ability
We use the GPT-2 architecture for pretraining, with large (1.5B) and small (137M) size models, using 8 GPUs (RTX A5000). Following [4], we set a max token length of 512, a global batch size 256, and a learning rate 2e-4, and train for 40K global steps (about 6B token duration). According to Ankner et al. [4], the relative performance trends observed at 40K steps are maintained in later training steps.
As our study requires a significant amount of resources for pre-training across multiple baselines, we carefully adopted settings and observations from recent representative works to ensure experimental rigor under limited resources. For example, [44, 4, 22] empirically shows that DSIR and PPL-based filtering mostly perform better than other rule-based (e.g., filtering based on character length, inclusion of certain words, and the ratio of English content) and model-based (e.g., FastText, memorization score, classifier on LLM embeddings) methods. We therefore regard DSIR and PPL-based filtering as representing the most advanced methods to date. Further explanation is in §A and §C.2.
@@
We provide further analyses and ablations for key issues in §B, such as preserving minority data, and sensitivity to factors (e.g., tokenizer, block size).”
@@ 4.2 Experiment on symbolic language corpus
Results The observed results are as follows: (1) Consistent with the previous experiments, the prior-based method achieves the best performance with significantly less time than the PPL-based approach. (2) These findings suggest that our methods holds not only for natural languages (e.g., English, Chinese) but also for artificial symbolic languages (e.g., code, math). This means that well-formed data in a certain language type can be identified via prior-based statistics, regardless of language type. (3) Math-related benchmarks (BIG-bench elementary math QA, GSM8K, SVAMP) exhibit near-random performance across all baselines, likely because the Pile-github dataset consists predominantly of code scripts.
@@ Appendix A Related works
(2) With reference: A well-known method is DSIR [44], which constructs embeddings using either n-grams or FastText on a curated dataset, and evaluates the similarity of samples in the raw dataset to these reference embeddings. DSIR is proposed to advance FastText [18]; it trains the model to classify a curated dataset as either high quality or low quality.
@@ Appendix B Additional analysis
Criteria Tokenizer Threshold Source Average normalized accuracy World knowledge Commonsense reasoning Language understanding Symbolic problem solving Reading compre- hension
@@ B.2 Sensitivity
We further conducted an experiment to determine which tokenizer yields better performance, and the result is reported in Table 4. Filtering based on GPT-2 and LLaMA-3 tokenizers performs almost identically, while T5 shows performance degradation. This can be interpreted as follows: We propose three possible interpretations: (1) There may exist a threshold of optimal granularity, which lies between 32K and 50K. (2) Using the tokenizer paired with the model may offer stability.
@@ C.1 Scores for each benchmark
ARC easy BIG-bench wikidata TriviaQA COPA OBQA PIQA HellaSwag LAMBADA Winograd Winogrande
@@
BIG-bench algorithms BIG-bench dyck lan- guages BIG-bench elementary math QA BIG-bench operators GSM8K SVAMP LSAT-LR LSAT-RC SAT-English CoQA
@@ C.2.1 Training token duration
We pretrain our baselines for 6B token duration, repeating 2 times on 3B corpus. This is to inherit the plausibility of the setting in [4]. As shown in Figure 1 of [4], the performance ranking of the baselines observed at 5B tokens remains consistent and stable across later stages of training. This shows that performance gaps between baselines are largely established around 5B token duration. Our work leverages this observation. Additionally, [22] also demonstrates similar patterns, where the performance ranking of the baselines remains consistent from the early stage.

### v2

@@ Abstract
As large language models (LLMs) are pretrained on massive web corpora, careful selection of data becomes essential to ensure effective and efficient learning. While perplexity (PPL)-based filtering has demonstrated strong performance, it suffers from drawbacks: substantial time costs and inherent unreliability of the model when handling noisy or out-of-distribution samples. In this work, we propose a simple yet powerful alternative: a prior-based data filtering method that estimates token priors using corpus-level term frequency statistics, inspired by linguistic insights on word roles and lexical density. Our approach filters documents based on the mean and standard deviation of token priors, serving as a fast proxy to PPL while requiring no model inference. Despite its simplicity, the prior-based filter achieves the highest average performance across 20 downstream benchmarks, while reducing time cost by over 1000× compared to PPL-based filtering. We further demonstrate its applicability to symbolic languages such as code and math, and its dynamic adaptability to multilingual corpora without supervision. The code is available online (https://github.com/ybseo-ac/prior_filter).
@@ 1 Introduction
The prior-based filter exhibits intriguing and practical properties, which we demonstrate empirically. (1) The linguistic principles underlying the term-frequency hold not only for English but also for other natural languages (e.g., Chinese and French), even for symbolic languages (e.g., code, mathematics). (2) Only a small amount of Chinese text data mixed into an English corpus may be noise and models can not learn patterns from it; however, as its amount increases, it becomes learnable by models. The prior-based filter is capable of automatically capturing this transition of learnability.
@@ 3.4.1 Prior-based filter approximates PPL-based filter
Figure 3: Extreme outlier samples selected based on three criteria, ensuring that each sample comes from a distinct criterion: PPL, $\mu_{d}$, and $\sigma_{d}$. ✓indicates filtered out.
@@
3.4.2 $\mu_{\mathtt{d}}$ reflects learnability of minor language in multi-lingual setting
The prior mean value has the property of dynamically reflecting the learnability of a data cluster (i.e., language type), especially when multiple clusters with distinct characteristics are mixed. For example, consider a corpus primarily composed of English data with a small portion of Chinese data included. While the Chinese samples may contain meaningful content, if their quantity is too small, the model will fail to learn the pattern of language. In this case, Chinese data is no more than noise. However, once the volume of Chinese data increases sufficiently, the model becomes able to interpret the language, making it learnable and meaningful data.
@@ 3.4.3 Fast, scalable filtering using subsampled priors
Remarkably, the already minimal computation time of the prior-based filter can be further reduced. For a 6B-token corpus, the entire process takes about 35 minutes on 40 CPUs (Intel Xeon Silver 4210R @ 2.40GHz), which consists of two stages: assessing the token prior, and computing $\mu_{\mathtt{d}}$ and $\sigma_{\mathtt{d}}$. Among these, the most time-consuming step is the token prior assessing phase, which alone takes around 30 minutes.
@@ 4.1 Experiment on natural language corpus and general ability
We use the GPT-2 architecture for pretraining, with large (1.5B) and small (137M) size models, using 8 GPUs (RTX A5000). Following [4], we set a max token length of 512, a global batch size 256, and a learning rate 2e-4, and train for 40K global steps (about 6B token duration). According to [4], the relative performance trends observed at 40K steps are maintained in later training steps.
As our study requires a significant amount of resources for pre-training across multiple baselines, we carefully adopted settings and observations from recent representative works to ensure experimental rigor under limited resources. For example, [44, 4, 22] empirically shows that DSIR and PPL-based filtering mostly perform better than other rule-based (e.g., filtering based on character length, inclusion of certain words, and the ratio of English content) and model-based (e.g., FastText, memorization score, classifier on LM embeddings) methods. We therefore regard DSIR and PPL-based filtering as representing the most advanced methods to date. Further explanation is in §A and §C.2.
@@
We provide further analyses and ablations for key issues in §B, such as preserving minority data, and sensitivity to factors (e.g., tokenizer, block size).
@@ 4.2 Experiment on symbolic language corpus
Results The observed results are as follows: (1) Consistent with the previous experiments, the prior-based method achieves the best performance with significantly less time than the PPL-based approach. (2) These findings suggest that our methods hold not only for natural languages (e.g., English, Chinese) but also for artificial symbolic languages (e.g., code, math). This means that well-formed data in a certain language type can be identified via prior-based statistics, regardless of language type. (3) Math-related benchmarks (BIG-bench elementary math QA, GSM8K, SVAMP) exhibit near-random performance across all baselines, likely because the Pile-github dataset consists predominantly of code scripts.
@@ Appendix A Related works
(2) With reference: A well-known method is DSIR [44], which constructs embeddings using either n-grams or FastText embeddings on a curated dataset, and evaluates the similarity of samples in the raw dataset to these reference embeddings. DSIR is proposed to advance FastText [18]; it trains the model to classify a curated dataset as either high quality or low quality.
@@ Appendix B Additional analysis
Criteria Tokenizer Threshold Source
Average
normalized
accuracy
World
knowledge
Commonsense
reasoning
Language
understanding
Symbolic
problem
solving
Reading
compre-
hension
@@ B.2 Sensitivity
We further conducted an experiment to determine which tokenizer yields better performance, and the result is reported in Table 4. Filtering based on GPT-2 and LLaMA-3 tokenizers performs almost identically, while T5 shows performance degradation. This can be interpreted as follows: We propose three possible interpretations: (1) There may exist a threshold of optimal granularity, which lies between 32K and 50K. (2) Using the tokenizer paired with the model may offer stability. Nevertheless, as demonstrated, priors can be computed even with tokenizers mismatched to the model. If an optimal tokenizer exists, it can be freely adopted at any time.
@@ C.1 Scores for each benchmark
ARC easy
BIG-bench
wikidata
TriviaQA COPA OBQA PIQA HellaSwag LAMBADA
Winograd
Winogrande
@@
BIG-bench
algorithms
BIG-bench
dyck lan-
guages
BIG-bench
elementary
math QA
BIG-bench
operators
GSM8K SVAMP LSAT-LR LSAT-RC SAT-English CoQA
@@ C.2.1 Training token duration
We pretrain our baselines for 6B token duration, repeating 2 times on 3B corpus. This is to inherit the plausibility of the setting in [4]. As shown in Figure 1 of [4], the performance ranking of the baselines observed at 5B tokens remains consistent and stable across later stages of training. This shows that performance gaps between baselines are largely established around 5B token duration. Our work leverages this observation.
@@
Additionally, [22] also demonstrates similar patterns, where the performance ranking of the baselines remains consistent from the early stage.
@@ Appendix D Details on Benchmarks
Appendix E Usage of AI assistants
In preparing this manuscript, we relied on AI-powered writing tools to refine sentence flow, fix grammatical mistakes, and improve readability. These assistants were used strictly for language polishing and played no role in shaping the technical content, research design, or experimental work. All scientific concepts, findings, and conclusions presented in this paper were fully developed and written by the researchers. The involvement of AI was limited to editorial support and did not influence the originality or intellectual contributions of the study.
Appendix F Filtered cases
In this section, we provide more cases that are classified as extreme outliers by the prior-based criteria.
F.1 Prior stds($\sigma_{\mathtt{d}}\$) based
Figure 7: Bottom-$n$ ranked data based on prior stds ($\sigma_{\mathtt{d}}$).
Figure 8: Top-$n$ ranked data based on prior stds ($\sigma_{\mathtt{d}}$). For the top 1–5 examples, the whitespace contains special characters other than “/s”.
d
d
F.2 Prior mean($\mu_{\mathtt{d}}\$)-based
Figure 9: Bottom-$n$ ranked data based on prior mean ($\mu_{\mathtt{d}}$).
Figure 10: Top-$n$ ranked data based on prior mean ($\mu_{\mathtt{d}}$).